OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

Parth (parthkohli):

\(x + x + 1 + x + 2 = 267\) Solve for x + 2

OpenStudy (anonymous):

how would i solve it? im not good at intgers

OpenStudy (lgbasallote):

let x be the first integer x + 1 second integer x+2 third integer sum of the three is 267 so x + x + 1 + x+ 2 = 267 3x + 3 = 267 3x = 267 - 3 3x = 264 x = 88 x + 1 = 89 x + 2 = 90

Parth (parthkohli):

Basically, three consecutive numbers can be represented by a number, a number more than that, and a number more than the second number.

OpenStudy (anonymous):

so is the answer 2?

Parth (parthkohli):

Solve the equation for x and then add two to x.

OpenStudy (anonymous):

im still confused...

OpenStudy (anonymous):

3x=264

OpenStudy (anonymous):

x + x +x = 267 - 2 - 1

OpenStudy (anonymous):

thilet us consider that digit is x then sum of three digits will x+x+1+x+2=267 3x+3=267 3x=267-3 3x=264 x=88 this is 1st digit 2nd digit was x+1 so 88+1=89 3rd was x+2 so 88+2=90 largest is 90...

OpenStudy (anonymous):

but it is consecutive integers . so its x+0 , x+1, x+2